OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = -32 and a9 = -120? Choose one answer. a. -384 b. -373 c. -362 d. -351

OpenStudy (anonymous):

Does anybody know how to do this?

OpenStudy (cwrw238):

a9 = a1 + 8d = -120 -32 + 8d = -120 8d = -88 d = -11 common difference = -11 32nd term = a1 + 31d = -32 + 31*-11 = -32 - 341 = -373

OpenStudy (anonymous):

Thanks!

OpenStudy (cwrw238):

ok - i was a bit slow as i was disturbed by an affectionate cat

OpenStudy (anonymous):

Haha its ok lol